# Advanced Video Inpainting Using Optical Flow-Guided Efficient Diffusion

### v2

@@ Abstract
Recently, diffusion-based methods have achieved great improvements in the video inpainting task. However, these methods still face many challenges, such as maintaining temporal consistency and the time-consuming issue. This paper proposes an advanced video inpainting framework using optical Flow-guided Efficient Diffusion, called FloED. Specifically, FloED employs a dual-branch architecture, where a flow branch first restores corrupted flow and a multi-scale flow adapter provides motion guidance to the main inpainting branch. Additionally, a training-free latent interpolation method is proposed to accelerate the multi-step denoising process using flow warping. Further introducing a flow attention cache mechanism, FLoED efficiently reduces the computational cost brought by incorporating optical flow. Comprehensive experiments in both background restoration and object removal tasks demonstrate that FloED outperforms state-of-the-art methods from the perspective of both performance and efficiency.
@@ 1 Introduction
Video inpainting aims to predict corrupted regions by filling them with contextually appropriate and temporally coherent content, which plays an essential role in computer vision, particularly in background restoration (BR) [23, 17, 34] and object removal (OR) [4, 33, 19].
Although conventional transformer-based approaches [15, 5, 37] have made great progress on standard benchmarks [31, 20], they encounter difficulties when requiring the synthesis of new content not present in existing frames. Currently, diffusion models [36, 38] have demonstrated significant prowess in text-guided video inpainting. However, temporal consistency remains an area with substantial scope for further improvement. And current methods struggle to generate satisfying results on BR and OR tasks, which is probably caused by the training priors of powerful base models. As demonstrated in Fig. 1, contents inpainted by CoCoCo [38] lacks harmony with its surroundings. This disharmony can be alleviated by providing motion guidance that is aligned with the surroundings. Optical flow, a key modality for motion information, provides significant guidance and enhances temporal consistency. Notably, integrating optical flow into video inpainting requires additional operations, including flow completion and effective incorporation, which incur extra computational costs. And diffusion inherently suffers from efficiency due to the multi-step denoising process. Consequently, when leveraging optical flow, it is essential to consider efficiency enhancements tailored to the multi-step characteristics of diffusion models.
Figure 1: Superiority of FloED on background restoration and object removal.
@@
Based on the above analysis, we propose an advanced video inpainting framework using optical Flow-guided Efficient Diffusion, called FloED. By leveraging motion information, our approach aims to enhance the reliability and performance of diffusion-based techniques in BR and OR applications. Initially, built on the large-scale T2I model, we inflate the origin layers and integrate pre-trained motion module [7] to establish a basic text-to-video framework. Then we fine-tune the motion module to effectively adapt it for the video inpainting domain. By adopting this framework as the primary inpainting branch, our innovations focus on three key aspects: (1) We design a tailored flow branch that completes corrupted flow while maintaining consistent channel numbers with the primary branch. Next, we integrate a multi-scale flow adapter that incorporates flow features into the decoder blocks of the primary U-Net architecture, enabling FloED to utilize motion information more effectively. (2) Building on the observation that adjacent latents share similar motion patterns [35] and recognizing that the essence of diffusion models involves fundamentally multi-step sampling processes, we introduce a training-free latent interpolation technique. This method leverages warping operation guided by optical flow to effectively accelerate the multi-step denoising process during early denoising stage. Furthermore, by incorporating a flow attention cache mechanism during the remain denoising stage as a complementary speed-up solution, we efficiently minimize the additional computational burden typically introduced by flow adapters and flow branches. (3) Recognizing that state-of-the-art image inpainting models significantly outperform video inpainting diffusion models, we utilize an anchor frame strategy to enhance the quality of video inpainting outcomes.
@@
  • We propose a dedicated dual-branch architecture that incorporates motion guidance with a multi-scale flow adapter, thereby enhancing temporal consistency and overall quality.
  • We introduce a training-free latent interpolation technique that leverages optical flow to speed up the multi-step denoising process. Complemented by a flow attention cache mechanism, FloED efficiently reduces the additional computational costs introduced by the flow.
  • We conducted extensive experiments, including both quantitative and qualitative evaluations, to validate that FloED demonstrates superiority over other state-of-the-art methods in both BR and OR tasks.
@@ 2 Related Work
2.1 Conventional Video Inpainting
Conventional video inpainting mainly focuses on two types of tasks: object removal and background restoration. Object removal [4, 33, 19] aims to eliminate unwanted objects from video frames. And background restoration [23, 17, 34] involves seamlessly reconstructing missing region of the background with coherent content.
Given the fact that it’s more simple to complete flows instead of directly filling masked regions [32], conventional transformer-based models [15, 35, 37] primarily rely on optical flow to propagate features or pixels for better inpainting outcomes. Specifically, E2FGVI [15] introduces an end-to-end framework with flow-guided feature propagation. FGT [35] combines decoupled spatiotemporal attention with a flow-guided content propagation. ProPainter [37] advances this field by merging dual-domain propagation with a mask-guided transformer.
Although propainter [37] has shown significant improvements on standard benchmarks [20, 31], it still struggles with generating coherent new content that isn’t present in existing frames. There is a pressing need to develop more robust solutions.
2.2 Diffusion-based Video Inpainting
In recent years, diffusion models [22, 9] have revolutionized the field of content generation, showcasing exceptional abilities in creating highly realistic outputs [21, 2]. The development of text-to-video (T2V) [7, 10] and image-to-video (I2V) [30, 3] diffusion generative models, has paved the way for innovations in text-guided video inpainting. Models like AVID [36] and CoCoCo [38], have highlighted the routine practice of employing pre-trained Stable Diffusion Inpainting models [2] equipped with motion modules [7] for text-guided video inpainting. Specifically, AVID [36] introduce zero-shot generation pipeline with a structure guidance module. CoCoCo [38] implements enhanced attention mechanisms to address text-video alignment and ensure motion consistency across frames. While these advancements have contributed to progress, existing diffusion-based video inpainting methods still exhibit significant room for improving temporal consistency and they perform suboptimally on both OR and BR tasks. Compared with them, FloED showcases distinct advantages over temporal consistency and overall quality.
Figure 2: Overview of FloED. FloED employs a dual-branch architecture implemented through a two-stage training approach. In the first training stage, we focus exclusively on the upper branch, optimizing the motion layer to adapt specifically to the video inpainting domain. Subsequently, we introduce a dedicated flow branch complemented by a multi-scale flow adapter, which provides flow guidance covering upblocks of primary UNet. During the inference phase, we enhance efficiency by integrating the flow attention cache (right part).
@@ 3 Preliminaries
where $\epsilon_{\theta}(\mathbf{z}_{t},t,c)$ is the predicted noise based on $\mathbf{z}_{t}$, the time step $t$ and the condition $c$. Once trained, we could leverage the deterministic sampling of DDIM [9] and PNDM [16] to denoise $\mathbf{z}_{t}$:
@@ 4 Methods
Building upon Stable Diffusion inpainting model [1], we integrate the motion module from Animatediff [7], adopting it as primary inpainting branch of our framework. As illustrated in Fig. 2, the training process of FloED can be divided into two stages. Initially, we adapt primary branch to video inpainting by fine-tuning the motion modules (Sec. 4.1). Our approach then incorporates a dedicated flow branch to complete corrupted flows estimated from masked frames, alongside a multi-scale flow adapter that provides motion guidance for the primary inpainting branch (Sec. 4.2). To further enhance video inpainting results, we implement an anchor frame strategy to leverage the priority of the image inpainting diffusion model [2] (Sec. 4.3). Furthermore, we introduce a training-free denoising acceleration technique that leverages optical flow for latent interpolation, compensated with a flow attention caching mechanism in the reference phase. We substantially enhance efficiency while significantly reducing the additional computational overhead the flow introduces. (Sec. 4.4).
4.1 Domain Adaptation
Given an original video sequence $\mathbf{V_{0}}=\{\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{N-1}\}\in\mathbb{R}^{N\times 3\times H\times W}$ and a binary mask sequence $\mathbf{m}=\{\mathbf{m}_{0},\mathbf{m}_{1},\ldots,\mathbf{m}_{N-1}\}\in\mathbb{R}^{N\times 1\times H\times W}$, corrupted frames $\mathbf{V_{m}}$ are obtained by applying the Hadamard product as follows: $\mathbf{V_{m}}=\mathbf{V_{0}}\odot\mathbf{m}$. We aim to generate a set of spatiotemporally consistent inpainted outcomes with temporally coherent and contextually appropriate content in the corrupted area while ensuring out-of-mask areas are unchanged.
Following previous solutions [36, 38], FloED is built on Stable Diffusion (SD) inpainting model [1], equipped with motion module from AnimateDiff-v3 [7] for better temporal consistency. Notably, during the first stage, only the primary branch is taken into consideration, and we only optimize the motion layer for adapting to the video inpainting.
Figure 3: Illustration of flow-guided latent interpolation (left) and warping operation (right) during the denosing process.
@@ 4.2 Flow Guidance in Video Inpainting
Typically, for video inpainting methods [15, 5, 37] that utilize optical flow, it is typically necessary to first complete the corrupted flow because only an estimated flow from the masked frames is initially available. Specifically, we first employ a pre-trained flow estimator [24] to perform corrupted flow estimation from $\mathbf{V_{m}}$. Subsequently, we introduce a flow sub-branch composed of the first layer from each block of the primary branch, which is independent of the timestep. As illustrated in Fig. 2, the flow sub-branch completes the corrupted flow while maintaining consistent channel numbers with the latent representations of the corresponding frames in the primary branch. This approach enables more effective utilization of motion information by integrating multi-scale flow adapters into all upblocks of the primary UNet. As for the flow adapter, it consists of learnable modules with decoupled cross-attention mechanisms that embed flow features into the primary branch (Fig. 2), thereby providing robust motion guidance effectively. We choose to insert the flow adapter between the text attention and the motion layer because it can compensate for the information provided by the prompt and further supply temporal guidance to the motion layer. Consequently, in the second training phase, we also trained the motion layer simultaneously. An additional optical loss $\mathcal{L}_{\textrm{flow}}$ is introduced in the second stage: ($F$ represents the ground truth flow.)
@@ 4.3 Anchor Frame Strategy
Recognizing that state-of-the-art image inpainting models significantly outperform their video inpainting counterparts, we introduce an anchor frame strategy to enhance the quality of video inpainting results. As illustrated in Fig. 2, for a given video sequence $\mathbf{V_{0}}$, we select an additional frame from the beginning of the sequence to serve as an anchor frame. We then utilize a pre-trained text-to-image (T2I) inpainting model [2] to reconstruct its corrupted region in advance. Subsequently, we concatenate the inpainted anchor frame with the noised video frames $\mathbf{V_{t}}$. This approach provides additional texture guidance to the video frames during the denoising process. After denoising, the anchor frame is discarded. This strategy leverages the superior performance of image inpainting models to improve the overall quality of video inpainting.
Figure 4: Qualitative Results. The orange dotted line indicates the input source.
@@ 4.4 Efficient Inference for FloED
Optical Flow Attention Cache Unlike the primary branch, the flow sub-branch is independent of timestep. During the inference phase, we utilize the flow sub-branch exclusively for flow completion in the first step and then use these completed flows for all subsequent steps. Regarding the flow adapter, it introduces additional computations by calculating the flow attention at every denoising step and multiple resolutions. To optimize this process, we establish a cache mechanism by computing the keys and values only during the first step and storing them in the memory bank (right part of Fig. 2). For the remaining steps, the cached keys and values are directly retrieved from the memory bank, eliminating the need for repeated calculations and enhancing efficiency.
Training-Free Denoising Speed-up Since adjacent feature latent exhibit similar motion patterns [35] and diffusion models generate high-level content early in the denoising process [13], we aim to speed-up the denoising process by interpolating latent using the completed flow. Notably, this technique is entirely training-free.
Specifically, as illustrated in Fig. 3, the initial step involves performing the standard denoising process for completing flow and caching flow attention. Subsequently, starting from step $t-1$, the noisy latent $\mathbf{z}$ is divided into two subsets based on parity. Latent interpolation process then follows a two-step alternating loop: even-indexed latents (shown in red) undergo denoising, while odd-indexed latents (shown in green) are obtained by warping operations using bi-directional optical flows. In the next step, only the interpolated latents (green) are denoised, and the red latents are generated through a similar warping process. Due to the negligible time cost of warping latent, the latency of denoising is halved by processing only half of the frame latent at each sampling timestep. Notably, the warping operation needs to be conducted at the $\mathbf{z}_{0}$ stage (as per Equation 4).
Figure 5: Qualitative comparisons. We compare our method against several approaches on BR and OR tasks.
Ideally, this process could be iterated until the final denoising step. However, since we are operating in latent space, completed optical flow provides only coarse-grained guidance, corresponding to the early stage of the diffusion process. Therefore, we restrict latent interpolation to the initial $S$ denoising steps, during which the overall structure of the image is established [13]. Additionally, to minimize flow errors, we perform warp operations exclusively between adjacent frames. Furthermore, to mitigate potential occlusion issues in flow warping [11], we perform the copy-paste operation in each denoising step (see Algorithm in appendix).

### v3

# Coherent Video Inpainting Using Optical Flow-Guided Efficient Diffusion

@@ Abstract
The text-guided video inpainting technique has significantly improved the performance of content generation applications. A recent family for these improvements uses diffusion models, which have become essential for achieving high-quality video inpainting results, yet they still face performance bottlenecks in temporal consistency and computational efficiency. This motivates us to propose a new video inpainting framework using optical Flow-guided Efficient Diffusion (FloED) for higher video coherence. Specifically, FloED employs a dual-branch architecture, where the time-agnostic flow branch restores corrupted flow first, and the multi-scale flow adapters provide motion guidance to the main inpainting branch. Besides, a training-free latent interpolation method is proposed to accelerate the multi-step denoising process using flow warping. With the flow attention cache mechanism, FLoED efficiently reduces the computational cost of incorporating optical flow. Extensive experiments on background restoration and object removal tasks show that FloED outperforms state-of-the-art diffusion-based methods in both quality and efficiency. Our codes and models will be made publicly available.
@@ 1 Introduction
Text-guided video inpainting aims to predict corrupted regions with text-aligned and temporally coherent contents, which has drawn widespread attention and applications. Conventional transformer-based methods [4, 17, 15, 31, 34] fundamentally lack textually semantic controllability when users require precise semantic alignment between textual descriptions and generated video results.
In recent years, diffusion models [21, 9, 20, 19] have demonstrated extraordinary capability in generating realistic and text-aligned content. Thus, diffusion-based solutions [10, 28, 35] have made substantial progress in the realm of text-guided image inpainting. However, directly applying these methods to video inpainting falls short in maintaining the necessary temporal consistency. Although rapid developments in text-to-video (T2V) generative diffusion models [6, 27] have successfully facilitated text-guided video inpainting [33, 36, 14] by leveraging motion modules [6] for temporal consistency, it still remains an area with substantial scope for further improvement. Meanwhile, current diffusion-based solutions struggle to produce satisfactory results in both Background Restoration (BR) and Object Removal (OR) scenarios. Both BR and OR need to generate background content semantically aligned with textual descriptions, while ensuring spatial-temporal coherence between synthesized content and contextual video semantics.
As evidenced in Fig. LABEL:fig:headfig, state-of-the-art methods like CoCoCo [36] exhibit spatial-temporal disharmony where inpainted regions demonstrate incompatible texture patterns or lighting conditions with surrounding contexts. We further hypothesize that this disharmony can be mitigated by integrating motion guidance that aligns with the scene. In this regard, optical flow, a key modality for capturing motion information, provides valuable guidance and enhances temporal consistency, potentially alleviating the observed inconsistencies. Notably, integrating optical flow into video inpainting requires additional operations, including flow estimation and flow completion and effective incorporation, all of which incur extra computational costs. Thus, diffusion models inherently suffer from efficiency due to the multi-step denoising process. Consequently, when we leverage optical flow, it is essential to consider efficiency enhancements tailored to the multi-step characteristics of diffusion models.
@@
Based on the above analysis, we propose a coherent video inpainting framework using optical Flow-guided Efficient Diffusion, called FloED. By leveraging motion information, our approach seeks to enhance both the performance and efficiency of diffusion-based techniques in BR and OR applications. Based on the Animatediff [6], we fine-tune the motion module in the first stage to effectively align its temporal modeling capacity with the video inpainting task. Built on this primary inpainting branch, our innovations focus on three key aspects: (1) We design a time-agnostic flow branch that completes corrupted flow while maintaining consistent channel numbers with the primary branch. Next, we integrate the multi-scale flow adapters that inject flow features into the decoder blocks of the primary U-Net architecture, enabling FloED to utilize motion information more effectively. (2) Building on the observation that adjacent latent features share similar motion patterns [31] and diffusion models fundamentally involve multi-step sampling processes, we introduce a training-free latent interpolation technique, which leverages warping operation guided by optical flow to effectively accelerate the multi-step denoising process during early denoising stage. Furthermore, by incorporating the flow attention cache mechanism during the remain denoising stage as a complementary speed-up solution, we efficiently minimize the additional computational burden typically introduced by flow adapters and the flow branch. (3) Recognizing that state-of-the-art image inpainting models significantly outperform video inpainting diffusion models, we utilize an anchor frame strategy to enhance the quality of video inpainting outcomes.
@@
  • Novel video inpainting model. We propose a dedicated dual-branch architecture that integrates optical flow guidance through flow adapters, thereby enhancing spatial-temporal consistency with compatible outcomes.
  • Efficient denoising process. We introduce a training-free latent interpolation technique that leverages optical flow to speed up the multi-step denoising process. Complemented by the flow attention cache mechanism, FloED efficiently reduces the additional computational costs introduced by the flow.
  • State-of-the-art performance. We conducted extensive experiments on OR and BR tasks, including both quantitative and qualitative evaluations, to validate that FloED outperforms other state-of-the-art text-guided diffusion methods in terms of both performance and efficiency.
Figure 1: Overview of FloED. FloED employs a dual-branch architecture implemented through a two-stage training approach. In the first training stage, we focus exclusively on the upper branch, optimizing the motion layer to adapt specifically to the video inpainting domain. Subsequently, we introduce a time-agnostic flow branch complemented by a multi-scale flow adapter, which provides flow guidance covering upblocks of primary UNet. During the inference phase, we enhance efficiency by integrating the flow attention cache (right part).
@@ 2 Related Work
Text-guided video inpainting, an emerging topic for synthesizing semantically consistent and temporally stable content in corrupted video regions through natural language instructions, has garnered significant research interest [33, 36, 14] and immense potential for creative and practical industrial applications. Although conventional transformer-based solutions [15, 31, 34] have shown significant improvements on standard benchmarks DAVIS [18] and Youtube-vos [29], they can’t provide semantic controllability when users require precise semantic alignment between textual descriptions and generated video results.
In recent years, diffusion models [21, 9] have revolutionized the field of content generation, showcasing exceptional abilities in creating realistic outputs [20, 19] and controllable outcomes [30, 32]. The application of diffusion models to text-guided image inpainting has seen remarkable advancements [10, 28, 35]. However, directly applying these methods to video inpainting falls short in maintaining the necessary temporal consistency. The emergence of text-to-video technologies [6, 27, 3], has provided the solutions for the text-guided video inpainting. Notably, Animatediff [6] preserves temporal coherence in video generation through motion module training while effectively utilizing Stable Diffusion’s robust framework [1]. Building upon this foundation, recent approaches [33, 36, 14] have advanced text-guided video inpainting. Specifically, AVID [33] proposes a fine-grained generation pipeline with structure guidance. CoCoCo [36] employs enhanced attention mechanisms to improve text-video alignment and motion consistency. And DiffuEraser [14] introduces the prior model for better inpainting outcomes. Nevertheless, existing methods demonstrate suboptimal performance on both object removal (OR) and background restoration (BR) tasks, frequently producing incompatible texture patterns or illumination conditions with surrounding contexts. In comparison, FloED exhibits superior temporal consistency and overall output quality.
@@ 3 Preliminaries
where $\epsilon_{\theta}(\mathbf{z}_{t},t,c)$ is the predicted noise based on $\mathbf{z}_{t}$, the time step $t$ and the condition $c$. Once trained, we could leverage the deterministic sampling of DDIM [9] to denoise $\mathbf{z}_{t}$:
@@ 4 Methods
Figure 2: Illustration of flow-guided latent interpolation (left) and warping operation (right) during the denoising process.
Given a text prompt $P$ and original video sequence $\mathbf{V_{0}}=\{\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{N-1}\}\in\mathbb{R}^{N\times 3\times H\times W}$ and a binary mask sequence $\mathbf{m}=\{\mathbf{m}_{0},\mathbf{m}_{1},\ldots,\mathbf{m}_{N-1}\}\in\mathbb{R}^{N\times 1\times H\times W}$, corrupted frames $\mathbf{V_{m}}$ are obtained by applying the Hadamard product as follows: $\mathbf{V_{m}}=\mathbf{V_{0}}\odot\mathbf{m}$. We aim to generate a set of spatiotemporally consistent inpainted outcomes with text-aligned contents which demonstrate compatible texture patterns and lighting conditions with surrounding context.
4.1 Network Overview
An overview of our proposed model, FLoED, is depicted in Fig. 1. For the architecture, FloED adopt pretrained Stable Diffusion Inpainting backbone [1] as the primary branch, while incorporating motion modules initialized from AnimateDiff v3 [6]. The training process of FloED can be divided into two stages.
In the first stage, we fine-tune the motion modules to align their temporal modeling capacity with video inpainting. In the second stage, FloED incorporates a dedicated flow branch to complete corrupted flows estimated from masked frames, alongside multi-scale flow adapters that inject hierarchical motion information into the primary inpainting branch (Sec. 4.2). To further enhance video inpainting results, we implement an anchor frame strategy to leverage the priority of the image inpainting diffusion model [2] (Sec. 4.3). Furthermore, we introduce a training-free denoising acceleration technique that leverages optical flow for latent interpolation, compensated with a flow attention caching mechanism in the reference phase. We substantially enhance efficiency while significantly reducing the additional computational overhead the flow introduces. (Sec. 4.4).
Notably, noised video latent $\mathbf{V_{t}}$, down-sampled binary mask $\mathbf{m}$, corrupted video latent $\mathbf{V_{m}}$ are concatenated along the channel axis as input.
@@ 4.2 Flow Guidance in Video inpainting
Considering the flow guidance, FloED first employs a pre-trained flow generator RAFT [22] to obtain the corrupted flow from $\mathbf{V_{m}}$, which is annotated as ”Corrupted Flow Estimation” in Fig. 1.
Flow Completion Branch.
To reconstruct the corrupted flow, the framework incorporates a dedicated flow completion branch that architecturally aligns with the primary inpainting backbone. Specifically, it selectively aggregate the initial ResNet module [7] from each corresponding block in the primary branch, to ensure channel-wise compatibility. And we exclude the time-step from ResNet to obtain the time-agnostic flow completion branch which creates stable flow feature regardless of diffusion progression. As shown in Fig 1, this flow completion branch enables comprehensive motion guidance by injecting compatible motion information into up-blocks of the primary UNet branch with multi-scale flow adapters.
Flow Adapter.
Our flow adapter is inspired by IP-Adapter [30], which is consisted with a separated cross-attention layer. The reconstructed flow features are fed into the cross attention for motion guidance. Notably, FloED strategically positions the flow adapter between the text cross-attention and motion modules to enable flow-aware latent modulation. This critical design further addresses misalignment caused by text-driven generation, because the multi-scale flow adapters dynamically adjusts latent features from the text cross-attention layer using optical flow priors, ensuring synthesized content maintains spatial-temporal compatibility with surrounding contexts. And in the second stage, we also continually fine-tune the motion modules.
The reconstructed optical loss $\mathcal{L}_{\textrm{flow}}$ is introduced in the second stage. ($F$ represents the ground truth flow)
@@ 4.3 Anchor Frame Strategy
Recognizing that state-of-the-art image inpainting models developed in academia far surpass their video inpainting counterparts, we introduce an anchor frame strategy that leverages these advanced image techniques to substantially improve video inpainting performance. As illustrated in Fig. 1, for a given video sequence $\mathbf{V_{0}}$, we select an additional frame from the beginning of the sequence to serve as an anchor frame. We then utilize a pre-trained text-to-image (T2I) inpainting model [2] to reconstruct its corrupted region in advance. Subsequently, we concatenate the inpainted anchor frame with the noised video frames $\mathbf{V_{t}}$. This approach provides additional texture guidance to the video frames during the denoising process. After denoising, the anchor frame is discarded. This strategy leverages the superior performance of image inpainting models to improve the overall quality of video inpainting.
@@ 4.4 Efficient Inference for FloED
Flow Attention Cache.
Unlike the primary branch, the flow branch is independent of timestep. During the inference phase, we utilize the flow branch exclusively for flow completion in the first step and then use these completed flows for all subsequent steps. Regarding the multi-scale flow adapters, they introduce additional computations by calculating the flow attention at every denoising step and multiple resolutions. To optimize this process, we establish a cache mechanism by computing the keys and values only during the first step and storing them in the memory bank (right part of Fig. 1). For the remaining steps, the cached keys and values are directly retrieved from the memory bank, eliminating the need for repeated calculations and enhancing efficiency.
Training-Free Denoising Speed-up.
Since adjacent feature latent exhibit similar motion patterns [31] and diffusion models generate high-level content early in the denoising process [12], we aim to speed-up the denoising process by interpolating latent using the completed flow. Notably, this technique is entirely training-free.
Specifically, as illustrated in Fig. 2, the initial step involves performing the standard denoising process for completing flow and caching flow attention. Subsequently, starting from step $t-1$, the noisy latent $\mathbf{z}$ is divided into two subsets based on parity. The latent interpolation process then follows a two-step alternating loop: even-indexed latents (shown in red) undergo denoising, while odd-indexed latents (shown in green) are obtained by warping operations using bi-directional optical flows. In the next step, only the interpolated latents (green) are denoised, and the red latents are generated through a similar warping process. Due to the negligible time cost of warping latent, the latency of denoising is halved by processing only half of the frame latent at each sampling timestep. Notably, the warping operation needs to be conducted at $\mathbf{z}_{0}$ (as per Equation 4).
Figure 3: Qualitative comparisons. We compare FloED against diffusion-based SOTAs on BR and OR tasks.
Ideally, this process could be iterated until the final denoising step. However, since we are operating in latent space, completed optical flow provides only coarse-grained guidance, corresponding to the early stage of the diffusion process. Therefore, we restrict latent interpolation to the initial $S$ denoising steps, during which the overall structure of the image is established [12]. Additionally, to minimize flow errors, we perform warp operations exclusively between adjacent frames. Furthermore, to mitigate potential occlusion issues in flow warping, we perform the copy-paste operation in each denoising step (see in appendix).